OpenStudy (anonymous):

Solve:What value of x makes the equation true? Answer(S) 7 seven-thirds equals x over six 10 14 21 Pick One Thanks For The Help (:

OpenStudy (anonymous):

it's 14, the reason being is if you multiply the bottom half of the fraction 3 x 2 is 6 so plug in the top with 2 and you get 14.

jimthompson5910 (jim_thompson5910):

7/3 = x/6 7*6 = 3x 42 = 3x 42/3 = x 14 = x x = 14

OpenStudy (anonymous):

you can do it both ways they both come out to the same answer

OpenStudy (anonymous):

Hope this helps :)

OpenStudy (anonymous):

Thanksss (Mud116 (: -

OpenStudy (anonymous):

no prob ;)